payton:

y = 4x – 10 y = 2 What is the solution to the system of equations? (3, 2) (2, 3) (–2, 2) (2, –2)

payton:

this ones easy isnt it?

ramen:

try figuring out this one on ur own. Ill give you the set up. 4x-10=2

ramen:

Yes! a lot easier

payton:

4x+12?

ramen:

with this one, the equations both have Y= to start, but with those that don't start with Y= for example 2x+y=4, we would plug the other Y= equation into Y. So it would be 2x+(Y=mx+b)=4

ramen:

YEs 4x+12 and i bet you know what to do after this

payton:

divide them together

ramen:

yes. 12/4

payton:

i got 3

payton:

and a is my only option with 3 as x

ramen:

yes good job!

ramen:

you are getting good at this

payton:

im trying as good as i can

ramen:

Ik you don't need to find Y bc u already know the answer, but in the future, to find Y you plug 3 into the x variable. So 2=4(3)-10

ramen:

but anyway you are doing great!